### v1

# RLeXplore: Accelerating Research in Intrinsically-Motivated Reinforcement Learning

Abstract
Extrinsic rewards can effectively guide reinforcement learning (RL) agents in specific tasks. However, extrinsic rewards frequently fall short in complex environments due to the significant human effort needed for their design and annotation. This limitation underscores the necessity for intrinsic rewards, which offer auxiliary and dense signals and can enable agents to learn in an unsupervised manner. Although various intrinsic reward formulations have been proposed, their implementation and optimization details are insufficiently explored and lack standardization, thereby hindering research progress. To address this gap, we introduce RLeXplore, a unified, highly modularized, and plug-and-play framework offering reliable implementations of eight state-of-the-art intrinsic reward algorithms. Furthermore, we conduct an in-depth study that identifies critical implementation details and establishes well-justified standard practices in intrinsically-motivated RL. The source code for RLeXplore is available at https://github.com/RLE-Foundation/RLeXplore.
11footnotetext: Equal Contribution.
1 Introduction
Reinforcement learning (RL) provides a framework for training agents to solve tasks by learning from interactions with an environment. At the core of RL is the optimization of a reward function, where agents aim to maximize cumulative rewards over time [1].
Figure 1: Episode Return achieved by the intrinsic rewards in RLeXplore in SuperMarioBros.
However, in complex environments, defining extrinsic rewards that effectively guide an agent’s learning process can be impractical, often requiring domain-specific expertise. In practice, poorly defined extrinsic rewards can lead to sparse-reward settings, where RL agents struggle due to the lack of a meaningful learning signal [2]. As the RL community tackles increasingly complex problems, such as training generally capable RL agents, there is a need for more autonomous agents capable of learning valuable behaviors without relying on dense supervision [3]. To address this challenge, the concept of intrinsic rewards has emerged as a promising approach in the RL community [4, 5, 6, 7, 8, 9]. Intrinsic rewards provide agents with additional learning signals, enabling them to explore and acquire skills across diverse environments beyond what extrinsic rewards alone can offer.
However, computing intrinsic rewards often requires learning auxiliary models, heavy engineering and performing expensive computations, making reproducibility challenging. While several formulations of intrinsic rewards have been proposed [5, 7, 10], each with its potential benefits for improving agent learning, the field lacks a comprehensive understanding of the comparative advantages and challenges posed by these formulations. Importantly, existing literature reports varying performance when using the same intrinsic rewards, reinforcing the need for a standardized framework and a deeper understanding of the optimization and implementation details.
In this paper, we introduce RLeXplore, an open-source library containing high-quality implementations of state-of-the-art intrinsic rewards. Our work presents a systematic study aimed at addressing gaps in understanding the critical implementation and optimization details of intrinsic rewards. To guide our investigation, we formulate numerous questions, aiming to uncover the intricacies of intrinsic rewards and their impact on RL agent performance. Our results highlight the importance of thoughtful implementation design for intrinsic rewards, showing that naive implementations can lead to suboptimal performance. Through carefully studied design decisions, we demonstrate significant performance gains.
Our contributions are threefold. Firstly, we provide a high-quality open-source repository for training RL agents, featuring the implementation of the most widely recognized intrinsic rewards. Secondly, we present a systematic evaluation that identifies the key implementation and optimization details critical to the success of intrinsic reward methods in RL. Lastly, we provide a comparative analysis of the agents’ performance across challenging environments, establishing a foundation for future research in intrinsically-motivated RL.
2 Related Work
Intrinsic rewards often rely on heavily engineered implementations to stabilize their optimization. The latter difficulties reproducibility and causes different works in the literature to report varying performance in popular environments. Some works have benchmarked intrinsic rewards in specific environments [11, 12, 10], yet they do not provide details on the importance of the design decisions in the implementation and optimization of the intrinsic rewards. In this work, we introduce RLeXplore, a more comprehensive framework that contains the most widely-used intrinsic rewards and provides the RL community with a unified framework to accelerate research and compare baselines in intrinsically-motivated RL. In the following, we overview existing formulations for intrinsic rewards of different natures and introduce the methods included in RLeXplore.
2.1 Count-Based Exploration
Count-based exploration methods provide intrinsic rewards by measuring the novelty of states, defined to be inversely proportional to the state visitation counts [13, 14, 15, 16]. In finite state spaces, count-based methods perform near optimally [13]. For this reason, these methods have been established as appealing techniques for driving structured exploration in RL. However, they do not scale well to high-dimensional state spaces [17, 18]. Pseudo-counts provide a framework to generalize count-based methods to high-dimensional and partially observed environments [17, 19, 20]. [4] proposed random network distillation (RND), which uses the prediction error against a fixed network as a learning signal that is correlated to counts. Recently, [8] used an elliptical bonus (E3B) and showed that such an objective provides a generalization of counts to high-dimensional spaces. In RLeXplore, we include Pseudo-counts, RND, and E3B as representatives of the state-of-the-art count-based methods.
2.2 Curiosity-Driven Exploration
Curiosity-based objectives train agents to interact with the environment seeking to experience outcomes that are not aligned with the agents’ predictions [21]. Hence, curiosity-driven exploration usually involves training an agent to increase its knowledge about the environment (e.g., environment dynamics) [22, 5, 23]. The intrinsic curiosity module (ICM) [5, 2] learns a joint embedding space with inverse and forward dynamics losses and was the first curiosity-based method successfully applied to deep RL settings. Disagreement [9] further extended ICM by using the variance over an ensemble of forward-dynamics models to compute curiosity. However, curiosity-driven methods are consistently found to be unsuccessful when the environment has irreducible noise [24]. To address the problem, [6] proposed RIDE, which uses the difference between two consecutive state embeddings as the intrinsic reward and encourages the agent to choose actions that result in significant state changes. In general, curiosity-based objectives remain amongst the most popular intrinsic rewards in deep RL applications to this day. In RLeXplore, we include ICM, Disagreement, and RIDE as representatives of the state-of-the-art curiosity-driven methods.
2.3 Global and Episodic Exploration
Towards more general and adaptive agents, recent works have studied decision-making problems in contextual Markov decision processes (MDPs) (e.g., procedurally-generated environments) [6, 8, 25]. Contextual MDPs require episodic-level exploration, where novelty estimates are reset at the beginning of each episode. [26] showed that both global and episodic exploration modalities have unique benefits and proposed combined objectives that achieve remarkable performance across many MDPs of different structures. NGU [7] and RIDE [6] also instantiate both global and episodic bonuses. Inspired by this recent line of works, in this paper, we study novel combinations of objectives for exploration that achieve impressive results in contextual MDPs.
3 Background
We frame the RL problem considering a MDP [27, 28] defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},E,P,d_{0},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $E:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the extrinsic reward function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition function that defines a probability distribution over $\mathcal{S}$, $d_{0}\in\Delta(\mathcal{S})$ is the distribution of the initial observation $\bm{s}_{0}$, and $\gamma\in[0,1)$ is a discount factor. The goal of RL is to learn a policy $\pi_{\bm{\theta}}(\bm{a}|\bm{s})$ to maximize the expected discounted return:
$J_{\pi}(\bm{\theta})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}E_{t}\right].$ (1)
Intrinsic rewards augment the learning objective to improve exploration. Letting $I:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denote the intrinsic reward function, the augmented optimization objective is:
$J_{\pi}(\bm{\theta})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(E_{t}+\beta_{t}\cdot I_{t})\right],$ (2)
where $\beta_{t}=\beta_{0}(1-\kappa)^{t}$ controls the degree of exploration, and $\kappa$ is a decay rate.
4 RLeXplore
In this section, we present RLeXplore, a unified, highly-modularized and plug-and-play framework that currently provides high-quality and reliable implementations of eight state-of-the-art intrinsic reward algorithms***RLeXplore complies with the MIT License.. Comparing multiple intrinsic reward methods under fair conditions is challenging due to various confounding factors, such as using distinct backbone RL algorithms (e.g., PPO [29], DQN [30], IMPALA [31]), optimization (e.g., reward and observation normalization, network architecture) and evaluation details (e.g., environment configuration, algorithm hyperparameters). RLeXplore is designed to provide a unified framework with standardized procedures for implementing, computing, and optimizing intrinsic rewards.
Figure 2: The workflow of RLeXplore. (a) RLeXplore monitors the agent-environment interactions and obtains data samples via a .watch() function. (b) For the sampled transitions, RLeXplore calculates the corresponding intrinsic rewards via a .compute() function and performs model updating via a .update() function.
4.1 Architecture
The core design decision of RLeXplore involves decoupling the intrinsic reward modules from the RL optimization algorithms, which enables our intrinsic reward implementations to be integrated with any desired RL algorithm (or existing library, see Appendix C). Figure 2 illustrates the basic workflow of RLeXplore, which consists of two parts: data collection (i.e., policy rollout) and reward computation.
At each time step, the agent receives observations from the environment and predicts actions. The environment then executes the actions and returns feedback to the agent, which consists of a next observation, a reward, and a terminal signal. During the data collection process, the .watch() function is used to monitor the agent-environment interactions. For instance, E3B [8] updates an estimate of an ellipsoid in an embedding space after observing every state. At the end of the data collection rollouts, .compute() computes the corresponding intrinsic rewards. Note that .compute() is only called once per rollout using batched operations, which makes RLeXplore a highly efficient framework. Additionally, RLeXplore provides several utilities for reward and observation normalization. Finally, the .update() function is called immediately after .compute() to update the reward module if necessary (e.g., train the forward dynamics models in Disagreement [9] or the predictor network in RND [4]). Section C illustrates the usage of the aforementioned functions. All operations are subject to the standard workflow of the Gymnasium API [32].
We provide various examples of using RLeXplore with popular RL frameworks (e.g., stable-baselines3 [33] and CleanRL [34]) in Appendix C.
4.2 Algorithmic Baselines
In RLeXplore, we implement eight widely-recognized intrinsic reward algorithms spanning the different categories described in Section 2, namely ICM [5], RND [4], Disagreement [9], NGU [7], PseudoCounts [7], RIDE [6], RE3 [35], and E3B [8], respectively. We selected them based on the following tenet:
  • The algorithm represents a unique design philosophy;
  • The algorithm achieved superior performance on well-recognized benchmarks;
  • The algorithm can adapt to arbitrary tasks and can be combined with arbitrary RL algorithms.
For detailed descriptions of each method we refer the reader to Appendix A.
5 Experiments
Figure 3: Screenshots of four selected exploration games, from left to right: Montezuma’s Revenge (MR), SuperMarioBros, Procgen-Maze, and Ant-UMaze.
We design our experiments with two objectives in mind: (i) evaluating the effectiveness of our implementations and their adaptability to different RL algorithms, and (ii) investigating the tuning of intrinsic rewards to provide well-justified standard implementations.
For the first part, we use the Montezuma’s Revenge (MR) and Ant-UMaze environments, shown in Figure 3. MR is a challenging exploration task from the Atari benchmark [36], while Ant-UMaze is an exploration and locomotion task from the Gymnasium robotics benchmark [37]. We provide the experimental settings and results in Appendix E and Appendix F. In these experiments, we seamlessly integrate RLeXplore with different frameworks (CleanRL [34], RLLTE [38]) and RL algorithms (PPO [29], SAC [39]).
For the second part, we use SuperMarioBros to investigate the low-level implementation details in intrinsic reward methods that lead to more robust performance. Additionally, we use Procgen-Maze to study how to best combine intrnisic and extrinsic rewards in sparse-reward tasks.
5.1 Low-level Implementation Details of Intrinsic Rewards
The performance of intrinsic rewards is affected by various factors, which tends to vary significantly with the complexity of the task, the RL algorithm used, the architecture of the networks, algorithm-specific hyperparameters, and the joint optimization of intrinsic and extrinsic rewards. As a result, implementing and reproducing intrinsic reward algorithms is challenging. To tackle this problem, we first formulate five research questions (RQs) to investigate how various low-level implementation details impact the training of intrinsically-motivated agents. We first define an initial baseline configuration for optimizing the intrinsic rewards, shown in Table 1. These baseline settings are selected based on the most common configurations reported in the literature. Next, we address each RQ sequentially, modifying the baseline configuration for each intrinsic reward as we gather new evidence regarding their critical implementation details. This iterative process leads to the development of high-quality implementations of state-of-the-art intrinsic reward methods.
In this section, we conduct reward-free experiments (i.e., without access to extrinsic rewards) using the SuperMarioBros environment [40]. SuperMarioBros is a widely used benchmark for evaluating exploration in RL [5, 6], as efficient exploration is closely related to effectively navigating the game levels and ultimately solving the game. Additional experimental settings can be found in Appendix B.
RQ 1: The impact of observation normalization.
Figure 4: Results for RQ 1, RQ 2, RQ 3, RQ 4, and RQ 5 in SuperMarioBros-1-1. The x-axis represents the episode return normalized by the performance of the extrinsic agent after training. A normalized return of 1.0 indicates that the agent solves the 1-1 level. Each bar represents the mean and standard deviation of the normalized episode returns. Results are aggregated over 5 seeds, and each run uses 10M environment interactions.
Observation normalization is crucial in deep learning to avoid numerical instabilities during optimization. Image observations, which typically range from 0 to 255, are commonly normalized to a range of 0 to 1 using Min-Max normalization by dividing by 255. However, previous studies suggest that Min-Max normalization may not be ideal for all representation learning algorithms [4].
In RQ 1, we compare Min-Max normalization with using an exponential moving average (EMA) of the mean and standard deviation for observation normalization (RMS). Figure 7 indicates that using RMS for observation normalization generally reduces the variance of the agent’s performance. Importantly, some intrinsic rewards, such as RND, NGU, PseudoCounts, and RIDE, critically require RMS normalization. This results indicate that RMS normalization is crucial for intrinsic reward methods that use random networks, since the lack of normalization can result in the embeddings produced by the random networks carrying very little information about the inputs [4].
RQ 2: The impact of reward normalization.
Similarly to RQ 1, reward normalization is also critical when using deep neural networks to compute the intrinsic rewards, since the scale of these rewards can be arbitrary and vary significantly over time. To mitigate the non-stationarity of intrinsic rewards, in RQ 2, we compare three normalization approaches: (1) Min-Max normalization, (2) using an RMS of the standard deviation, and (3) no reward normalization.
Surprisingly, our findings reveal that not normalizing intrinsic rewards at all can improve the performance of some intrinsic reward methods (e.g., RE3 and ICM). However, our results show that Min-Max reward normalization is a considerable option, improving the performance of the majority of the methods.
RQ 3: The co-learning dynamics of policies and representations for intrinsic rewards.
Optimizing intrinsic rewards in deep RL often involves training representations in auxiliary networks (e.g., predictor network in RND, inverse dynamics encoder in ICM, forward dynamics encoders in Disagreement). However, managing the co-learning dynamics of representations and policies is challenging. In RQ 3, we explore three update strategies for representations: (1) updating representations at the same frequency as the policy, (2) updating representations 50% of the time, and (3) updating representations 10% of the time. This comparison sheds light on the trade-off between the number of gradient updates in the representations and the performance of the policy.
Our findings suggest that, in general, the learned representations of the auxiliary networks need to be caught up (i.e., equally trained) with the same data as the RL agents. Our results show that training the representations slowly (e.g., using only 10% of the data at each iteration) can cause significant performance decreases for global intrinsic rewards like ICM, Disagreement and RND. In contrast, for intrinsic rewards that utilizes episodic memories (e.g., NGU and RIDE), training the representations slowly may improve the performance to a certain extent. These observations indicate that training intrinsic rewards and RL agents does not always have to be synchronized entirely. Moreover, we can reduce the computational overhead by reducing the training time of intrinsic rewards.
RQ 4: The impact of weight initialization.
Weight initialization plays a crucial role in optimizing deep neural networks, enabling faster convergence. In RQ 4, we compare two approaches for weight initialization in the auxiliary networks: (1) orthogonal weight initialization and (2) uniform weight initialization (PyTorch’s default). Note that the policy and value networks remain unchanged.
Our results show the importance of weight initialization in intrinsically-motivated RL. Specifically, we observe that using orthogonal initialization of weights leads to notable performance improvements for most intrinsic rewards, regardless of their specific optimization tasks (e.g., inverse dynamics, forward dynamics), and even in random networks (e.g., RND and RE3).
RQ 5: Is memory required to optimize intrinsic rewards?
In RQ 5, we investigate whether the intrinsic rewards included in RLeXplore benefit from memory-enabled architectures. We compare the optimization of intrinsic rewards using a vanilla network and one equipped with a long-short term memory (LSTM) [41] module, while keeping PPO as the RL backbone algorithm.
Some intrinsic reward methods exhibit lower performance when using LSTM policies. This observation aligns with the fact that LSTMs provide episodic context to policies, whereas most intrinsic reward methods define exploration as a global problem. Interestingly, for RIDE, which computes the state embedding changes as the intrinsic rewards, the episodic context provided by LSTMs enables agents to better optimize the intrinsic reward.
Figure 5: Overall (left) and individual (right) performance comparisons between the baselines and best configurations. For the left figure, interquartile mean (IQM), mean, median, and optimality gap (OG, lower is better) are computed over all the algorithms and runs. PoI is the probability of improvement of the best configurations compared to the baselines. Our results show that carefully-tuned implementations for the intrinsic rewards achieve significantly better exploration and task performance.
Based on the results from RQ 1-5, we have identified the optimal configurations that surpass the predefined baselines. Reward and observation normalization are notably influential on performance, and by adjusting the update proportion and using an LSTM, it is possible to further boost the performance of intrinsic rewards. Figure 5 demonstrates both the comprehensive and individual performance comparisons before and after tuning. The average probability of improvement across all intrinsic rewards is 82%, with performance increases of 86% and 66% for the aggregate IQM and mean, respectively.
5.2 Combination of Intrinsic and Extrinsic Rewards
RQ 6: Joint Optimization of Intrinsic and Extrinsic Rewards
Training agents to maximize two learning signals concurrently can be challenging. In sparse-reward environments, the objective is for agents to explore the state space by optimizing intrinsic rewards until they discover the task rewards, at which point they should focus solely on optimizing the task rewards. However, many intrinsically-motivated RL applications naively optimize the sum of intrinsic and extrinsic rewards, potentially leading to learning fuzzy value functions and suboptimal policies. In this section, we compare this approach with learning two separate value functions, one for each stream of rewards. The advantages of the latter include the ability to disentangle the effects of intrinsic and extrinsic rewards on the agent’s behavior, leading to clearer learning dynamics and potentially more efficient exploration.
For this analysis, we used the Procgen-Maze task [42] as a sparse-reward benchmark. RL agents often struggle to learn meaningful behaviors from the extrinsic reward alone in this task. We evaluate different variants of the task (e.g., 1 maze vs. 200 mazes) to examine singleton versus contextual MDPs. Figure 6 demonstrates that learning two separate value functions, which we refer as the TwoHead architecture, outperforms the naive approach of simply adding the two rewards in the complex sparse-reward environment of Procgen-Maze, both in singleton and contextual settings. Importantly, all methods outperform the extrinsic agent, especially in the 1 Maze environment.
Figure 6: (Left) During training, the extrinsic agent struggles to find the goal in the selected Maze, resulting in a reward of 0. While some intrinsic reward methods yield occasional non-zero rewards, the algorithms perform significantly better when intrinsic and extrinsic value estimation are decoupled using two distinct value heads in the agent’s network. (Right) In the Procgen variant where each maze represents a unique level, the baseline extrinsic agent achieves the goal 50% of the time, and intrinsic rewards don’t outperform the baseline significantly. We note that the presence of easier levels, where the goal may occasionally be near the agent’s starting point results in generally less sparse rewards and an easier task to learn.
5.3 Unlocking the Potential of Intrinsic Rewards
RQ 1-6 extensively discuss the tuning of intrinsic rewards under both normal and reward-free scenarios, revealing significant insights into the optimization processes. However, we aim to delve deeper into the capabilities of intrinsic rewards to address the evolving challenges in the RL community. Specifically, in RQ 7 and 8, we investigate recent developments in the exploration literature in RL, such as combined intrinsic rewards and exploration in contextual MDPs. For our experiments, we use the SuperMarioBrosRandomStages environment variant, where agents play a different level in the game at each episode. Our results indicate that the recent developments in combined intrinsic rewards merit further research, as we demonstrate that such methods can enable agents to learn exploratory behaviors of exceptional quality in both singleton and contextual MDPs.
RQ 7: Which intrinsic rewards generalize better in contextual MDPs?
In contextual MDPs, there is little shared structure across episodes, since the episodic context can vary the environment significantly. In this settings, global intrinsic rewards, which re-use experience from past episodes to compute novelty in the current episode can provide wrong estimations. Conversely, episodic intrinsic rewards, such as E3B and PseudoCounts, are specifically designed to estimate novelty within each new episode, aligning better with the dynamic nature of contextual MDPs. As shown in Figure 15, E3B achieves the highest performance among all the intrinsic rewards, while other intrinsic rewards struggle to adapt and nearly fail to learn. This distinct advantage underscores the importance of designing intrinsic rewards that are context-sensitive and capable of updating their novelty detection mechanisms based on the specific characteristics of each episode.
RQ 8: The performance of mixed intrinsic rewards.
Finally, we study the potential of combined intrinsic rewards [26]. We run experiments using all the levels in the game of SuperMarioBros, and we sample them uniformly during training. As in RQ 1-5, we do not use the extrinsic reward for training the agents but use it as an evaluation metric to show how much agents actively explore the environment.
Our results show that combined objectives enable emergent behaviors of much better quality than single objectives. Interestingly, E3B and RIDE are the best performing single objectives, and E3B+RIDE also achieves the highest performance among all the combinations. Similarly, RND and ICM, combined with other intrinsic rewards, outperform their original performance. This indicates that different intrinsic rewards can provide orthogonal gains that can be leveraged together.
Figure 7: (Left) The performance ranking of single and mixed intrinsic rewards on the SuperMarioBrosRandomLevels. As expected, episodic bonuses (such as E3B and RIDE) demonstrate superior performance, attributed to the environment’s non-singleton MDP nature. (Right) Overall performance comparisons between the single and mixed intrinsic rewards. Here, G.E. denotes the six "global+episodic" combinations, and G.G. denotes the three "global+global" combinations, as illustrated in Table 3.
6 Conclusion
As the RL community continues to tackle increasingly complex tasks, the adoption of intrinsic rewards promises to play a pivotal role in improving agent autonomy and adaptability in diverse environments. In conclusion, RL agents can benefit significantly from intrinsic rewards. These enable the discovery of emergent behaviours that can make RL agents more autonomous and proficient in solving complex sparse-reward tasks. Our work introduces RLeXplore, a comprehensive open-source repository that not only implements state-of-the-art intrinsic rewards but also provides a systematic evaluation framework for understanding their impact on agent performance. We design RLeXplore to accelerate research in intrinsically-motivated RL and support fair comparsisons with well-stablished baseliens.
Through our study, we emphasize the importance of thoughtful implementation design, demonstrating that well-considered approaches lead to significant performance gains over naive implementations. Our contributions extend to establishing standardized practices for implementing and optimizing intrinsic rewards, laying the groundwork for future advancements in intrinsically motivated RL.
Our results show that with RLeXplore, RL agents can learn emergent behaviours autonomously, solving multiple levels of SuperMarioBros without task rewards. Additionally, we show that intrinsic rewards enable RL agents to solve complex sparse-reward tasks like Procgen-Maze. Finally, RLeXplore facilitates further research in combined intrinsic rewards, uncovering the potential of such methods.
Appendix A Algorithmic Baselines
ICM [5]. ICM leverages a inverse-forward model to learn the dynamics of the environment and uses the prediction error as the curiosity reward. Specifically, the inverse model inferences the current action $\bm{a}_{t}$ based on the encoded states $\bm{e}_{t}$ and $\bm{e}_{t+1}$, where $\bm{e}=\psi(\bm{s})$ and $\psi(\cdot)$ is an embedding network. Meanwhile, the forward model $f$ predicts the encoded next-state $\bm{e}_{t}$ based on $(\bm{e}_{t},\bm{a}_{t})$. Finally, the intrinsic reward is defined as
$I_{t}=\|f(\bm{e}_{t},\bm{a}_{t})-\bm{e}_{t+1}\|_{2}^{2}.$ (3)
RND [4]. RND produces intrinsic rewards via a self-supervised manner, in which a predictor network $\hat{f}$ is trained to approximate a fixed and randomly-initialized target network $\hat{f}$. As a result, the agent is motivated to explore unseen parts of the state space. The intrinsic reward is defined as
$I_{t}=\|\hat{f}(\bm{s}_{t+1})-f(\bm{s}_{t+1})\|_{2}^{2}.$ (4)
Disagreement [9]. Disagreement is variant of ICM that leverages an ensemble of forward models and calculates the intrinsic reward as the variance among these models. Accordingly, the intrinsic reward is defined as
$I_{t}=\mathrm{Var}\{f_{i}(\bm{e}_{t},\bm{a}_{t})\},i=0,...,N$ (5)
NGU [7]. NGU is a mixed intrinsic reward approach that combines global and episodic exploration and the first algorithm to achieve non-zero rewards in the game of Pitfall! without using demonstrations or hand-crafted features. The intrinsic reward is defined as
$I_{t}=\min\{\max\{\alpha_{t}\},C\}/\sqrt{N_{\rm ep}(\bm{s}_{t})},$ (6)
where $\alpha_{t}$ is a life-long curiosity factor computed following the RND method, $C$ is a chosen maximum reward scaling, and $N_{\rm ep}$ is the episodic state visitation frequency computed by pseudo-counts.
PseudoCounts [7]. Pseudo-counts has been widely used in count-based exploration approaches [17, 19] with diverse implementations like neural density models. In this paper, we follow NGU [7] that computes pseudo-counts via $k$-nearest neighbor estimation, which is highly efficient and can be applied to arbitrary task. Given the encoded observations $\{\bm{e}_{0},\dots,\bm{e}_{T-1}\}$ visited in the an episode, we have
$\sqrt{N_{\rm ep}(\bm{s}_{t})}\approx\sqrt{\sum_{\tilde{\bm{e}}_{i}}K(\tilde{\bm{e}}_{i},\bm{e}_{t})}+c,$ (7)
where $\tilde{\bm{e}}_{i}$ is the first $k$ nearest neighbors of $\bm{e}$, $K$ is a Dirac delta function, and $c$ guarantees a minimum amount of pseudo-counts. Finally, the intrinsic reward is defined as
$I_{t}=1/\sqrt{N_{\rm ep}(\bm{s}_{t})}$ (8)
RIDE [6]. RIDE is designed based on ICM that learns the dynamics of the environment and rewards significant state changes. Accordingly, the intrinsic reward is defined as
$I_{t}=\|\bm{e}_{t+1}-\bm{e}_{t}\|_{2}/\sqrt{N_{\rm ep}(\bm{s}_{t+1})},$ (9)
where $N_{\rm ep}(\bm{s}_{t+1})$ is used to discount the intrinsic reward and prevent the agent from lingering in a sequence of states with a large difference in their embeddings.
RE3 [35]. RE3 is an information theory-based and computation-efficient exploration approach, which aims to maximize the Shannon entropy of the state visiting distribution. In particular, RE3 leverages a random and fixed neural network to encode the state space and employs a $k$-nearest neighbor estimator to estimate the entropy efficiently. Then the estimated entropy is transformed into particle-base intrinsic rewards. Specifically, the intrinsic reward is defined as
$I_{t}=\frac{1}{k}\sum_{i=1}^{k}\log(\|\bm{e}_{t}-\tilde{\bm{e}}_{t}^{i}\|_{2}+1).$ (10)
E3B [8]. E3B provides a generalization of count-based rewards to continuous spaces using an elliptical bonus. E3B learns a representation mapping from observations to a latent space (e.g., using inverse dynamics). At each episode, the sequence of latent observations parameterize an ellipsoid which is used to measure the novelty of the subsequent observations. In tabular settings, the E3B ellipsoid reduces to the table of inverse state-visitation frequencies [8]. The reward of a new observation is computed as:
$I_{t}=f(\bm{e}_{t})^{T}C_{t-1}f(\bm{e}_{t})$ (11)
where $f$ is the learned representation mapping and $C_{t-1}$ is the episodic ellipsoid.
Appendix B Experimental Settings
B.1 Baselines
We designed the following settings for the baseline experiments, and all the subsequent RQs were progressively adjusted based on the baselines. Moreover, all the experiments are performed using the proximal policy optimization (PPO) [29] implementation from RLLTE [38].
Table 1: Details of baseline settings.
Hyperparameter Value
Observation normalization image = image / 255.0
Reward normalization RMS
Weight initialization Orthogonal
Update proportion 1.0
with LSTM False
Table 2: PPO hyperparameters for SuperMarioBros and Procgen games.
Hyperparameter SuperMarioBros Procgen
Observation downsampling (84, 84) (64, 64)
Stacked frames No No
Environment steps 10000000 25000000
Episode steps 128 256
Number of workers 1 1
Environments per worker 8 64
Optimizer Adam Adam
Learning rate 2.5e-4 5e-4
GAE coefficient 0.95 0.95
Action entropy coefficient 0.01 0.01
Value loss coefficient 0.5 0.5
Value clip range 0.1 0.2
Max gradient norm 0.5 0.5
Epochs per rollout 4 3
Batch size 256 2048
Discount factor 0.99 0.999
B.2 Details of RQs
Table 3: Details of candidates for all RQs, where $\bm{\mathrm{I}}$ is a batch of intrinsic rewards.
# Candidate Detail
RQ 1 Min-Max image = image / 255.0
RMS $\mathrm{image}=\mathrm{Clip}\left(\frac{\mathrm{image}-\mathrm{running\>mean}}{\mathrm{running\>std.}},-5.0,5.0\right)$
RQ 2 Vanilla $\bm{\mathrm{I}}=\bm{\mathrm{I}}$
RMS $\bm{\mathrm{I}}=\frac{\bm{\mathrm{I}}}{\mathrm{running\>std}}$
Min-Max $\bm{\mathrm{I}}=\frac{\bm{\mathrm{I}}-\min(\bm{\mathrm{I}})}{\max(\bm{\mathrm{I}})-\min(\bm{\mathrm{I}})}$
RQ 3 0.1 Use 10% of the samples to update the intrinsic reward module.
0.5 Use 50% of the samples to update the intrinsic reward module.
1.0 Use 100% of the samples to update the intrinsic reward module.
RQ 4 Vanilla Fill the input tensor with values drawn from the uniform distribution.
Orthogonal Fill the input tensor with a (semi) orthogonal matrix.
RQ 5 Vanilla Policy network with only convolutional and linear layers.
LSTM Policy network that includes an LSTM layer.
RQ 6 Vanilla $R=E+I$
Two-head Value network uses two separate branches for $E$ and $I$.
RQ 7 N/A N/A
RQ 8 Global+Episodic
E3B+RND, E3B+ICM, E3B+RIDE,
RE3+RND, RE3+ICM, RE3+RIDE
Global+Global RND+ICM, RND+RIDE, ICM+RIDE
Table 3 illustrates the details of the candidates for all RQs. For RQ 1-5, we designed the experiments sequentially and modified the configuration for each intrinsic reward based on the best results of previous RQs. For instance, experiments of RQ 1 only change the technique of observation normalization, and RQ 2 will use the best observation normalization method for each reward module obtained in RQ 1. Likewise, RQ 3 will follow the best results obtained in RQ 1-2 and only change the proportion of samples used for model update. However, we kept using the baselines settings for each reward in RQ 8 to explore the most original performance of the mixed intrinsic rewards.
B.3 Best Configurations
Table 4: The best configurations for each intrinsic reward on SuperMarioBros games.
Reward Obs. Norm. Reward Norm. Update Prop. Weight Init. Memory Required
PseudoCounts Min-Max 0.5 Orthogonal
ICM RMS 1.0 Orthogonal
RND Vanilla 0.1 Orthogonal
E3B Min-Max 0.1 Orthogonal
RIDE Min-Max 0.1 Default
RE3 Vanilla N/A Orthogonal
NGU Min-Max 0.1 Orthogonal
Disagreement Min-Max 1.0 Orthogonal
Appendix C Usage Examples
C.1 Workflow of RLeXplore
[…]
C.2 RLeXplore with Stable-Baselines3
[…]
C.3 RLeXplore with CleanRL
[…]
Appendix D Learning Curves
Figure 8: Learning curves of the baselines and RQ 1 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 9: Learning curves of the RQ 2 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 10: Learning curves of the RQ 3 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 11: Learning curves of the RQ 4 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 12: Learning curves of the RQ 5 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 13: Learning curves of RQ 6 on Procgen-1MazeHard. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 14: Learning curves of RQ 6 on Procgen-AllMazeHard. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 15: Learning curves of RQ 7 on SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 16: Learning curves of RQ 8 (global+episodic exploration) on SuperMarioBros-1-1-v3 and SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 17: Learning curves of RQ 8 (global+global exploration) on SuperMarioBros-1-1-v3 and SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Appendix E On-Policy RL Algorithms and Discrete Control Tasks
In this section, we demonstrate the combination of RLeXplore and on-policy RL algorithms and its effectiveness on discrete control tasks. Specifically, we couple the PPO algorithm and intrinsic rewards, and evalulate their performance on Montezuma’s Revenge, a hard exploration task from the Atari benchmark [36]. Meanwhile, we utilize the PPO implementation of CleanRL [34] to show the adaptability of RLeXplore. Table 5 illustrates the training hyperparameters used for the experiments.
Table 5: Training hyperparameters for Montezuma’s Revenge.
Part Hyperparameter Value
Observation downsampling (84, 84)
Stacked frames 4
Environment steps 1e+8
Episode steps 128
Number of workers 1
Environments per worker 8
Optimizer Adam
PPO Learning rate 1e-4
GAE coefficient 0.95
Action entropy coefficient 0.01
Value loss coefficient 0.5
Value clip range 0.1
Max gradient norm 0.5
Epochs per rollout 4
Batch size 256
Discount factor 0.99
Observation normalization RMS
Reward normalization RMS
Intrinsic reward Weight initialization Orthogonal
Update proportion 0.25
with LSTM False
Figure 18: Since only RND can achieve significant results in this task among the eight intrinsic rewards, we only show the results of RND. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Appendix F Off-Policy RL Algorithms and Continuous Control Tasks
To showcase the generality of RLeXplore, we run additional experiments in settings different from the ones in the main paper. Concretely, we couple intrinsic rewards with soft actor-critic (SAC) [39], an off-policy RL algorithm, and test their performance in Ant-UMaze, a continuous control task with sparse rewards. Table 6 illustrates the training hyperparameters used for the experiments. We show the performance of Disagreement, RND, ICM and vanilla SAC in Figure 19. The results indicate that intrinsically-motivated agents are able to navigate the maze more efficiently, finding the goals more often than the vanilla agents that can only learn from the sparse task rewards.
We only use 3 intrinsic rewards with SAC because of the episodic nature of the other intrinsic reward methods. For example, the episodic memory in RIDE, PseudoCounts, NGU; and the episodic ellipsoid in E3B require the replay buffer to sample entire episodes instead of random rollouts. We aim to implement this logic in the future in our RLeXplore codebase.
Table 6: Training hyperparameters for Ant-Umaze.
Part Parameter Value
Total timesteps $1\cdot 10^{6}$
Buffer size $1\cdot 10^{6}$
Discount ($\gamma$) 0.99
Target smoothing coefficient ($\tau$) 0.005
Batch size 256
Learning starts 5000
Policy learning rate $3\cdot 10^{-4}$
Q function learning rate $1\cdot 10^{-3}$
Policy frequency 2
Target network frequency 1
Noise clip 0.5
Entropy coefficient ($\alpha$) 0.2
Auto-tune entropy coefficient True
Observation normalization RMS
Reward normalization RMS
Intrinsic reward Weight initialization Orthogonal
Update proportion 0.25
with LSTM False
Figure 19: Performance comparison between the three selected intrinsic rewards and the extrinsic reward.